OpenStudy (anonymous):

You work 2 days a week for 5 hours for $7.25. How much would you have in 1 month

OpenStudy (anonymous):

290$

OpenStudy (beccaboo333):

Please do not give direct answers.

OpenStudy (anonymous):

its not a question in any b ook or work sheet I needa know my pay at my job

OpenStudy (beccaboo333):

There are 4 weeks in a month, if you do 2x4 you get 8. You work 8 days a month. Now what's 8x5?

OpenStudy (beccaboo333):

Once you get that you would then times that by 7.25 to get your answer. Welcome to OpenStudy c:

OpenStudy (anonymous):

7.25*5=36.25 per day 36.25*2=72.5 per week 72.5*4=290 per month

OpenStudy (anonymous):

thank you......its still not enough to get the plane ticket so thanks for the help anyways

OpenStudy (beccaboo333):

Ask for a couple extra hours or days a week then. Good luck c:

OpenStudy (anonymous):

im stupid lol I thought we were in june..........how much would I makev in 2

OpenStudy (anonymous):

290x2

OpenStudy (anonymous):

would that 580 be june 2

OpenStudy (anonymous):

No, June 2nd is about a month away, so it would still be $290

OpenStudy (anonymous):

by the end of june like 25 would be? I needa get the ticket by then